OpenStudy (anonymous):

How do I solve the quadratic equation by completing the square? -3x^2 + 9x = 1

OpenStudy (anonymous):

I have the solution but that isn't one of my choices. x = 1/6(9 + sqrt69)

OpenStudy (anonymous):

Well first, in completing the square, we need to simplify the equation as much as possible. Let's factor out -3 from it. -3(x^2 - 3x) = 1

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

And set to 0. Where, -3(x^2 - 3x - 1) = 0

OpenStudy (anonymous):

Okay, whats next?

OpenStudy (anonymous):

im wondering if the end result will start with -3 or a fraction

OpenStudy (anonymous):

Actually I apologize very sincerely. This is how you do it.